OpenStudy (anonymous):

(x12)4 = x3 x8 x16 x48

jagr2713 (jagr2713):

first we Simplify (x^12)^4 For all positive integer exponents (a^n)^m = a^(m n). Apply this to (x^12)^4. Multiply exponents. (x^12)^4 = x^(12×4): x^(12×4) Multiply 12 and 4 together. 12×4 = 48:

OpenStudy (anonymous):

so basically i just multiply both the numbers and thats it @jagr2713